# Large Language Models Could Be Rote Learners

### v4

@@ Abstract
Multiple-choice question (MCQ) benchmarks are widely used for evaluating Large Language Models (LLMs), yet their reliability is undermined by benchmark contamination. In this study, we reframe contamination as an inherent aspect of learning and seek to disentangle genuine capability acquisition from superficial memorization in LLM evaluation. First, by analyzing model performance under different memorization conditions, we uncover a counterintuitive trend: LLMs perform worse on memorized MCQs than on non-memorized ones, indicating the coexistence of two distinct learning phenomena, i.e., rote memorization and genuine capability learning. To disentangle them, we propose TrinEval, a novel evaluation framework reformulating MCQs into an alternative trinity format, reducing memorization while preserving knowledge assessment. Experiments validate TrinEval’s effectiveness in reformulation, and its evaluation reveals that common LLMs may memorize by rote 20.5% of knowledge points (in MMLU on average).
footnotetext: ${}^{\text{\Letter}}$orresponding Author: Haochao Ying.
1 Introduction
Figure 1: MCQ-based evaluation. We observe that LLMs tend to underperform on memorized MCQs.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs), driven primarily by large-scale pre-training on massive datasets, has endowed these models with remarkable proficiency across diverse tasks (Ouyang et al., 2022; OpenAI, 2024; Touvron et al., 2023). As LLMs continue to improve, evaluating their genuine capacities has emerged as a fundamental challenge, necessitating proper methodologies to ensure fairness and robustness (Ganguli et al., 2023; Liu et al., 2023b).
@@
Among the developed methods, multiple-choice question (MCQ) benchmarks have become a standard approach for evaluation. Typically, LLMs are presented with a question and a fixed set of answer choices, requiring them to select the most appropriate option (see Fig. 1 for illustration). This format enables straightforward performance measurement through accuracy metrics and could cover a wide range of subjects. However, despite their widespread adoption, MCQ-based evaluation raises concerns about reliability due to benchmark contamination (Li and Flanigan, 2024; Kim et al., 2024), i.e., test data unintentionally appears in training corpora and models may exploit memorized content rather than demonstrating genuine understanding, inflating their apparent capabilities. For instance, Zhou et al. (2023) discovers that smaller models with deliberate pre-exposure could outperform their larger counterparts, thereby contradicting widely accepted scaling laws.
@@
To mitigate the issue, Zhou et al. (2023) advocates the removal of benchmark datasets from pre-training corpora. However, this strategy conflicts with the fundamental objective of large-scale pre-training, which aims to maximize model performance by exposing LLMs to as much data as possible. From a broader perspective, human learning also involves problem-solving through practicing on similar questions, e.g., exam preparation. While rote memorization of specific questions and answers merely lead to short-term success, repeated practicing can also facilitate deeper conceptual understanding. Inspired, rather than viewing benchmark contamination as a flaw to be eradicated, which is a nearly impossible task at scale (Sainz et al., 2023; Bordt et al., 2024), we argue that it is an inherent aspect of learning and should be accounted for in evaluation. Therefore, this study shifts its focus to evaluating LLMs in the presence of contamination, aiming to distinguish genuine capability gains from superficial memorization effects. The explicit disentangling of these two learning effects remains largely unexplored in MCQ-based evaluation, yet we believe it marks a crucial step towards developing more rigorous and unbiased evaluation methodologies.
@@
To investigate the effects of superficial memorization in LLM evaluation, we compare model performance under different memorization conditions. Inspired by membership inference attacks (MIA) (Carlini et al., 2022a, 2021), we define superficial memorization as an LLM’s ability to verbatim reproduce content, e.g., MCQs in our case. Using this criterion, we partition the MMLU benchmark (Hendrycks et al., 2020) [^1] Selected for its popularity and documented data contamination in widely used LLMs (Sainz et al., 2023). into memorized and non-memorized subsets and evaluate three open-source LLMs [^2] Llama2-7B (Touvron et al., 2023), Mistral-7B-v0.2 (Jiang et al., 2023) and Vicuna-v1.5-7B (Zheng et al., 2023b). on both. Surprisingly, results reveal a consistent yet counterintuitive trend: LLMs perform worse on memorized MCQs than on those not (see Fig. 1 for illustration and Fig. 2 for results). This challenges the assumption that memorization improves model performance and suggests the coexistence of two distinct learning phenomena in LLMs: rote memorization, where models recall content verbatim without true understanding, and genuine capability learning, where they internalize underlying knowledge.
The preliminary investigation has several limitations. First, the binary classification of MCQs as either memorized or non-memorized oversimplifies the nuances of memorization, potentially overlooking intermediate cases. Second, we rely on accuracy to measure performance, which is inherently unreliable. Third, our analysis could not reveal the mutual effects between rote memorization and capability learning. To address these challenges, we propose TrinEval, a novel evaluation framework designed to provide a more reliable measure of LLM performance by minimizing the influence of rote memorization. TrinEval employs a query-based probing (q-probing) mechanism Allen-Zhu and Li (2023) that reformulates MCQs into an alternative trinity format, i.e., entity-attribute-context. This could prevent direct content recall while preserving knowledge assessment.
@@
Through experiments, we demonstrate that TrinEval’s reformulation is knowledge-preserving, i.e., maintaining testing problems’ inherent knowledge requirements without introducing extra cues, and could effectively reduce memorization. Combined with a continuous superficial memorization quantification metric, TrinEval reveals the in-robustness of LLMs’ capability learning, e.g., with MMLU, tested open-sourced LLMs only mastered 19.6% of knowledge points while 20.5% are memorized by rote in the meanwhile, shedding light on the necessity for further optimization.
2 Related Work
2.1 LLM Evaluation on MCQ Benchmarks
The rapid advancement of LLMs has driven their expansion into diverse domains, necessitating robust and fair evaluation methodologies (Zheng et al., 2023b; Hu et al., 2025) and platforms (Contributors, 2023; Chiang et al., 2024). Among these, evaluating on MCQ benchmarks emerges as a widely adopted approach due to the ease of validation and standardized comparison across models (Hendrycks et al., 2020; Wang et al., 2024; Zhong et al., 2023; Huang et al., 2024).
However, MCQ-based evaluations are not without limitations. Biases in LLM responses have been extensively studied (Dai et al., 2024), revealing issues such as social biases (Salewski et al., 2024; Liu et al., 2023a) and order sensitivity (Akter et al., 2023). To mitigate the latter, Pride (Zheng et al., 2023a) estimates the option positional bias after option permutation. To examine mastery of knowledge, Zhao et al. (2023) applies a hypothesis testing method and checks rephrased-context consistency for a given question. Benchmark contamination is arguably the most severe challenge for MCQ-based evaluations, which may result in misleadingly inflated performance (Zhou et al., 2023; Li and Flanigan, 2024). To address this, prior studies have explored data filtering, frequently-updated test sets (White et al., 2025), and data perturbation (Li et al., 2024).
@@ 2.1 LLM Evaluation on MCQ Benchmarks
In this paper, instead of attempting to eliminate contamination, we evaluate LLMs under its presence, aiming to distinguish genuine capability gains from superficial memorization effects. This marks a new perspective of LLM evaluation, revealing the extent to which models truly understand concepts rather than merely memorizing data.
2.2 LLM Memorization
@@
Membership inference attacks (MIA) are commonly used to determine whether a specific sample was present in a model’s training data. Initially studied in smaller models, Carlini et al. (2022b) investigates deep learning memorization mechanisms by identifying and removing easily detectable memorized samples. In the context of LLMs, MIA has been employed to assess privacy risks, revealing that both open- and closed-source models can leak sensitive personal data when provided with related prompts (Kim et al., 2024).
@@
Beyond privacy concerns, Carlini et al. (2022a) formally defines LLM memorization as a model’s ability to verbatim generate text sequences following a prefix prompt. Using this definition, several studies (Sainz et al., 2023; Bordt et al., 2024; Carlini et al., 2021) have examined mainstream LLMs, confirming widespread test data leakage across popular benchmarks. To quantify memorization strength, researchers (Shi et al., 2023; Zhang et al., 2024; Oren et al., 2023; Carlini et al., 2019) have further explored methods such as analyzing token probability distributions in generated outputs. However, while these studies extensively analyze LLM memorization, few explicitly investigate how memorization influences an LLM’s problem-solving ability. In contrast, our work focuses on their interplay, presenting a more rigorous approach to fair and reliable LLM evaluation.
3 Methodology
3.1 Pre-investigation of LLM Capability w.r.t. Memorization
@@ 3 Methodology
Benchmark contamination often leads to inflated performance estimate. This phenomenon is commonly attributed to models memorizing specific questions and answers rather than demonstrating genuine problem-solving abilities. However, the extent to which and how memorization influences LLM performance remains unclear. To disentangle genuine capability acquisition from superficial memorization, we conduct a preliminary investigation into how LLMs perform under different memorization conditions. By examining model accuracy on memorized vs. non-memorized subsets, we aim to reveal the role of memorization in LLM evaluation and establish a foundation for more rigorous assessment methodologies.
Formally, we define an MCQ as $x=\{x_{Q},x_{O},x_{W}\}$, where $x_{Q}$, $x_{O}$, and $x_{W}$ refer to the question, options, and ground-truth answer, respectively. Following the memorization definition from Carlini et al. (2022a), we say an MCQ $x$ is memorized by LLM $G$ if $G$ can extract/generate the content of options $x_{O}$ exactly given question $x_{Q}$. In practice, we incorporate meta-information (e.g., benchmark name) and 5-shot examples to recall memory and use greedy decoding (i.e., temperature fixed to 0) during extraction (Bordt et al., 2024; Sainz et al., 2023) (refer to Appendix A for the complete prompt). Using MMLU (Hendrycks et al., 2020) as the evaluation benchmark, we divide the test set MCQs into memorized and non-memorized subsets, where the memorized subset consists of 909–982 questions (accounting for 6.5%–7.0% of the total 14,006) depending on the tested LLMs Llama2-7B, Mistral-7B-v0.2, and Vicuna-v1.5-7B. The detailed statistics of questions across subsets are given in Table 1 of Appendix A, and we also observe that the majority of memorized questions are those relatively simple, i.e., not in MMLU-PRO (Wang et al., 2024).
@@
We then compute the accuracy (ACC) of tested LLMs by subsets as a proxy of model performance under different memorization conditions. The results of both zero- and five-shot prompting are reported in Fig. 2, from which we observe a consistent yet somehow counterintuitive trend: LLMs exhibit 47.2% lower accuracy on average on memorized MCQs compared to non-memorized ones, regardless of LLMs and prompting techniques. This finding challenges the commonly held assumption that memorization directly improves model performance. In addition, it also implies the coexistence of two distinct learning paradigms within LLMs, which we term rote memorization and genuine capability learning, respectively.
@@
However, our pre-investigation has its limitations. The binary classification of memorization potentially overlooks more nuanced forms of learning. Additionally, using ACC as the performance metric does not truly capture model capacity. We address these two issues in the following subsections, which then ensure a disentangle analysis of rote memorization and capability learning.
3.2 Quantifying LLM Memorization
For quantifying the memorization of LLMs, prior research (Shi et al., 2023; Zhang et al., 2024) suggests that outlier tokens, which exhibit higher generation probabilities, are more likely to be found in memorized samples. Building on this idea, we develop a metric that utilizes the bottom $K\%$ of token probabilities within the generated sequence as a measure of memorization. Formally, the memorization score $F_{m}(\overline{x},G)$ of LLM $G$ on text sequence $\overline{x}$ is computed as follows:
where $p_{G}(\overline{x}_{i}|\overline{x}_{1:i-1})$ denotes the generation probability of token $\overline{x}_{i}$ by $G$ given its prefix subsequence as context, and set $\mathcal{M}_{K}(\overline{x})$ includes the $K\%$ of tokens with the lowest probabilities. The higher $F_{m}$ is, the more likely $\overline{x}$ is memorized by the LLM, i.e., the least memorized content could still been extracted with a high probability.
3.3 Measuring LLM Capability with TrinEval
We next present TrinEval, a novel evaluation framework designed to provide a more reliable measure of LLM performance by minimizing the influence of rote memorization.

### v5

@@ Abstract
Benchmark-based evaluation, e.g., multiple-choice questions (MCQs) and open-ended questions (OEQs), is widely used for evaluating Large Language Models (LLMs), yet their reliability is undermined by benchmark contamination. When pre-exposed to testing benchmarks during training, less capable LLMs have been found to achieve inflated performance, thereby yielding erroneous results. In this study, we reframe contamination as an inherent aspect of learning and seek to disentangle and expose genuine capability acquisition from superficial memorization in LLM evaluation. Following this, firstly, by analyzing model performance under different memorization conditions of MCQs, we uncover a counterintuitive trend: LLMs perform worse on memorized benchmarks than on non-memorized ones, indicating the coexistence of two learning phenomena, i.e., rote memorization and genuine capability learning. To disentangle them, we propose TrinEval, a novel evaluation framework that reformulates MCQs into an alternative knowledge-centric trinity format, reducing memorization while preserving inherent knowledge, enabling the evaluation of genuine capability in the presence of memorization. Extensive experiments validate the effectiveness and robustness of TrinEval in reformulating benchmarks, and the evaluation results further reveal that mainstream LLMs rely on rote memorization for an average of 19.6% of knowledge points across the MMLU and the GSM8K dataset.
I Introduction
@@
The rapid advancement of Large Language Models (LLMs), driven primarily by large-scale pre-training on massive datasets, has endowed these models with remarkable proficiency across diverse tasks [47, 13, 7]. As LLMs continue to improve, evaluating their genuine capacities has emerged as one of the fundamental challenges, necessitating proper methodologies to ensure fairness and robustness [48, 49].
@@
Among the developed methods, benchmark-based evaluation has become a standard approach to evaluate LLMs. Typically, taking multiple-choice questions (MCQs) as an example, LLMs are presented with a question and a fixed set of optional choices, requiring them to select the most appropriate answer (see Fig. 1 for illustration). This format enables straightforward performance measurement through accuracy metrics and could cover a wide range of subjects. However, despite their widespread adoption, benchmark-based evaluation raises concerns about reliability due to benchmark contamination [35, 36], i.e., test data unintentionally appears in training corpora, and LLMs may exploit memorized content rather than demonstrating genuine understanding, inflating their apparent capabilities. For instance, Zhou et al. [34] discovers that smaller models with deliberate pre-exposure could outperform larger counterparts, thereby contradicting widely accepted scaling laws.
Figure 1: Benchmark-based evaluation. We observe that LLMs tend to underperform on memorized questions.
@@
To mitigate this issue, Zhou et al. [34] advocates the removal of benchmarks from pre-training corpora. However, this strategy conflicts with the fundamental objective of large-scale pre-training, which aims to maximize model performance by exposing LLMs to as much data as possible. From a broader perspective, human learning also involves problem-solving through practicing on similar questions, e.g., exam preparation. While rote memorization of specific questions and answers merely leads to short-term success, repeated practice can also facilitate deeper conceptual understanding. Thus, rather than viewing benchmark contamination as a flaw to be eradicated, which is a nearly impossible task at scale [3, 4], we argue that it is an inherent aspect of learning and should be accounted for in evaluation. Therefore, this study shifts focus to evaluating LLMs in the presence of contamination, aiming to distinguish genuine capability gains from superficial memorization effects. The explicit disentangling of these two learning effects remains largely unexplored in benchmark-based (MCQ-based, as an example) evaluation, yet we believe it marks a crucial step towards developing more rigorous and unbiased evaluation methodologies.
@@
To investigate superficial memorization effects in LLM evaluation, we compare model performance under different memorization conditions. Inspired by membership inference attacks (MIA) [1, 37], we consider the intuitive manifestation of superficial memorization is the LLM’s ability to verbatim reproduce content, e.g., extracting options of MCQs in our case. With the criterion, we partition the MMLU benchmark [6] [^1] Selected for its popularity and documented data contamination in widely used LLMs [3]. into memorized and non-memorized subsets and evaluate three open-source LLMs [^2] Llama2-7B [7], Mistral-7B-v0.2 [11] and Vicuna-v1.5-7B [12].. Surprisingly, results reveal a consistent yet counterintuitive trend: LLMs perform worse on memorized MCQs than on those not (see Fig. 1 for illustration and Fig. 2 for results). This challenges the assumption that memorization inflates model performance and suggests the coexistence of two learning phenomena in LLMs: rote memorization, where models recall content verbatim without true understanding, and genuine capability learning, where they internalize underlying knowledge. Still, our preliminary investigation needs further exploration. First, binary classification of benchmarks as either memorized/non-memorized and capable/incapable oversimplifies the nuances of memorization and capability, potentially overlooking intermediate cases. Second, our analysis could not reveal the numerical dependency between rote memorization and capability learning.
@@
To address these challenges, we propose TrinEval, a novel evaluation framework designed to provide a reliable measure of LLM performance by minimizing the influence of rote memorization. TrinEval employs a query-based probing (q-probing) mechanism [18] that reformulates benchmarks into an alternative trinity format, i.e., entity-attribute-context, which prevent direct content recall while preserving knowledge assessment, and exposes the genuine capability of LLMs. Through experiments, we demonstrate that TrinEval’s reformulation is knowledge-preserving, i.e., maintaining testing problems’ inherent knowledge requirements without introducing extra cues, and could effectively reduce memorization. Combined with a continuous superficial memorization quantification metric, TrinEval reveals the in-robustness of LLMs’ capability in benchmark evaluation. With MMLU, tested open-sourced LLMs only mastered 19.0% of knowledge points while 19.6% are memorized by rote in the meanwhile, while for OEQs in GSM8K, LLMs only mastered 17.12% of knowledge points while 17.30% are rote memorized, shedding light on the necessity for further optimization on the LLMs development. Our main contributions are as follows:
  • We propose a novel evaluation paradigm that assesses the genuine capability of LLMs under the default condition of benchmark contamination.
  • We introduce TrinEval, a framework that gets access to the genuine knowledge stored within tokens through textual reformulation and explores LLMs’ true understanding based on verbal queries.
  • Through TrinEval’s memorization-capability decoupling and quantitative analysis, we reveal a counterintuitive but general finding: the more LLMs memorize, the worse their actual problem-solving capability becomes.
II Related Work
II-A LLM Evaluation on Benchmarks
The rapid advancement of LLMs has driven their expansion into diverse domains, necessitating robust and fair evaluation methodologies [12, 8] and platforms [24, 25]. Among these, evaluating on MCQ benchmarks emerges as a widely adopted approach due to the ease of validation and standardized comparison across models, including MMLU [6], a large-scale multi-task benchmark covering 57 subjects spanning STEM, humanities, social sciences, and professional disciplines, designed to assess general knowledge and reasoning ability; AGIEval [27], a human-centric benchmark constructed from real-world public examinations such as college entrance exams and certification tests, focusing on practical knowledge application and high-level cognitive skills; C-Eval [28], a comprehensive Chinese multi-level evaluation suite covering middle school, high school, university, and professional domains, tailored for evaluating foundation models in Chinese understanding and reasoning; and MMLU-Pro [5], an enhanced version of MMLU that filters out simplistic questions and retains complex, model-challenging items to better measure robust high-order reasoning capacity. However, MCQ-based evaluations are not without limitations. Biases in LLM responses have been extensively studied [26], revealing issues such as social biases [30, 29] and order sensitivity [31]. To mitigate the latter, PriDe [22] estimates the option positional bias after option permutation. To examine mastery of knowledge, Zhao et al. [32] apply a hypothesis testing method and check rephrased-context consistency for a given question. Benchmark contamination is arguably the most severe challenge for MCQ-based evaluations, which may result in misleadingly inflated performance [34, 35]. To address this, prior studies have explored data filtering, frequently-updated test sets [9], and data perturbation [10], yet none of these can fundamentally resolve the issue.
Beyond MCQs-based benchmarks, open-ended questions (OEQs) evaluation has gained increasing attention for assessing LLMs’ generative capabilities, reasoning chains, and factual consistency in free-form responses. Representative open-ended benchmarks include GSM8K [53], which focuses on multi-step mathematical reasoning with detailed chain-of-thought annotations; MATH [6], covering advanced algebraic, geometric, and combinatorial problems; TruthfulQA [57], designed to measure truthfulness and avoid falsehoods in generation; and CommonsenseQA 2.0 [59] for commonsense understanding in free-response settings. These benchmarks require LLMs to produce coherent, logically complete, and factually accurate answers rather than selecting from fixed options, making them more aligned with real-world applications. Still, OEQ-based evaluation faces similar challenges from benchmark contamination. Since test samples, including full questions, reasoning steps, and reference answers, are frequently leaked into training corpora, LLMs may directly reproduce memorized rationales instead of performing genuine reasoning [35, 3]. Also, evaluating OEQ responses introduces additional difficulties: automated metrics such as BLEU, ROUGE, or BERTScore often fail to capture logical correctness, while human evaluation is costly, inconsistent, and difficult to scale [8, 12]. Recent work has attempted to mitigate these issues by perturbing question templates [10], using dynamic test sets [9], or applying model-based judges [24, 8], yet few methods [60] explicitly disentangle rote memorization from genuine reasoning ability in open-ended generation and provide a commonly agreed evaluation metric. As a result, reliably evaluating the true problem-solving capacity of LLMs under contamination remains an open challenge for both MCQ and OEQ benchmarks.
@@ II-A LLM Evaluation on Benchmarks
In this paper, instead of attempting to eliminate contamination, we evaluate LLMs under its presence, aiming to distinguish genuine capability gains from superficial memorization effects. This marks a new perspective on LLM evaluation, revealing the extent to which models truly understand concepts rather than merely memorizing data.
II-B LLM Memorization
@@
Membership inference attacks (MIA) are commonly used to determine whether a specific sample was present in a model’s training data. Initially studied in smaller models, Carlini et al. [33] investigates deep learning memorization mechanisms by identifying and removing easily detectable memorized samples. In the context of LLMs, MIA has been employed to assess privacy risks, revealing that both open- and closed-source models can leak sensitive personal data when provided with related prompts [36].
@@
Beyond privacy concerns, Carlini et al. [1] formally defines LLM memorization as a model’s ability to verbatim generate text sequences following a prefix prompt. Using this definition, several studies [3, 4, 37] have examined mainstream LLMs, confirming widespread test data leakage across popular benchmarks. To quantify memorization strength, researchers [16, 17, 38, 39] have further explored methods such as analyzing token probability distributions in generated outputs. However, while these studies extensively analyze LLM memorization, few explicitly investigate how memorization influences an LLM’s problem-solving ability. In contrast, our work focuses on their interplay, presenting a more rigorous approach to fair and reliable LLM evaluation.
III Methodology
III-A Pre-investigation of LLM Capability w.r.t. Memorization
Benchmark contamination often leads to inflated performance estimate. This phenomenon is commonly attributed to models memorizing specific question and answer sequence rather than demonstrating genuine problem-solving abilities. However, the extent to which and how memorization influences LLM performance remains unclear. To disentangle genuine capability acquisition from superficial memorization, we conduct a preliminary investigation into how LLMs perform under different memorization conditions. By examining model accuracy on memorized vs. non-memorized subsets, we aim to reveal the role of memorization in LLM evaluation and establish a foundation for more rigorous assessment methodologies.
In this paper, we mainly discussed the LLM Capability w.r.t. Memorization via MCQs as an example [^3] We also discuss and validate the extension of TrinEval on OEQs in Sec. IV-D.. Formally, we define an MCQ as $x=\{x_{Q},x_{O},x_{W}\}$, where $x_{Q}$, $x_{O}$, and $x_{W}$ refer to the question, options, and ground-truth answer, respectively. Following the memorization definition from Carlini et al. [1], we define an MCQ $x$ is memorized by LLM $G$ if $G$ can extract the option contents $x_{O}$ exactly given the question $x_{Q}$. In practice, given the commonly used next-token prediction pre-training, we incorporate meta-information (e.g., benchmark name) and 5-shot examples to elicit the memorized following content  [^4] We aim to evoke the memorization of LLMs by prompting this information. For detailed reasons, please refer to Sec. IV-C., and use greedy decoding (i.e., temperature fixed to 0) during extraction [4, 3] (refer to Appendix A for the complete prompt). Using MMLU [6] as the evaluation benchmark, we divide the test set MCQs into memorized and non-memorized subsets, where the memorized subset consists of 909–982 questions (accounting for 6.5%–7.0% of the total 14,006) depending on the tested LLMs LlaMA2-7B, Mistral-7B-v0.2, and Vicuna-v1.5-7B. The detailed statistics of questions across the memorized/non-memorized subsets are given in Tab. 1 in the Appendix A.
@@ III Methodology
We then compute the accuracy (ACC) of tested LLMs by subsets as a proxy of model performance under different memorization conditions. The results of both zero- and five-shot prompting are reported in Fig. 2, from which we observe a consistent yet somehow counterintuitive trend: LLMs exhibit 47.2% lower accuracy on average on memorized MCQs compared to non-memorized ones, regardless of LLMs and prompting techniques. This finding challenges the commonly held assumption that memorization directly improves model performance. In addition, it also implies the coexistence of two distinct learning paradigms or extremes within LLMs, which we term rote memorization and genuine capability learning, respectively. We also observe that the majority of memorized questions are those relatively simple, i.e., not in MMLU-PRO [5] (refer to Tab. 1 in the Appendix A), which is a commonly recognized difficult subset of MMLU . Since MMLU-Pro’s curation prioritizes complex, model-challenging questions by excluding simplistic ones, the non-inclusion of our memorized questions in MMLU-Pro serves as objective evidence of their relative simplicity. This alignment with MMLU-Pro’s explicit filtering criterion validates our claim as reasonable, eliminating the possibility that the memorized quenstions are relatively difficult.
@@
However, our pre-investigation has its limitations. Using ACC as the performance metric does not truly capture model capacity. Additionally, the binary classification of memorization potentially overlooks more nuanced forms of learning due to the discrete classification cutoff. We address these two issues in the following Sec. III-B and Sec. III-C, respectively, which then ensure a disentangled and quantitative analysis of rote memorization and capability learning.
III-B Measuring LLM Capability with TrinEval
First, in order to truly capture model capacity, we present TrinEval, a novel evaluation framework designed to provide a reliable measure of LLM performance by minimizing the influence of rote memorization. TrinEval first preliminarily extract the knowledge-centric triplet of each question, detecting where the knowledge is stored and how to expose the knowledge through query-based probing. The framework then verify and reflect whether the trinity is qualified or not. Finally, TrinEval utilizes the trinity to evaluate the genuine capability of LLMs.